OpenStudy (dragon123mama):

Solve the equation 0.25R - 0.125 + 0.5R = 0.5 + R

OpenStudy (dragon123mama):

When i did this i only got is half right. i am not sure where i went wrong.

OpenStudy (p0sitr0n):

put constants on one side, all terms with Rs on the other one: 0.25R+0.5R-R=0.5+0.125 -0.25R=0.625 R=-2.5

OpenStudy (dragon123mama):

ohhh i see now. thanks :)